# Actor-Critic Model Predictive Control: Differentiable Optimization meets Reinforcement Learning

### v6

@@ Abstract
An open research question in robotics is how to combine the benefits of model-free reinforcement learning (RL)—known for its strong task performance and flexibility in optimizing general reward formulations—with the robustness and online replanning capabilities of model predictive control (MPC). This paper provides an answer by introducing a new framework called Actor-Critic Model Predictive Control. The key idea is to embed a differentiable MPC within an actor-critic RL framework. This integration allows for short-term predictive optimization of control actions through MPC, while leveraging RL for end-to-end learning and exploration over longer horizons. Through various ablation studies, we expose the benefits of the proposed approach: it achieves better out-of-distribution behaviour, better robustness to changes in the dynamics and improved sample efficiency. Additionally, we conduct an empirical analysis that reveals a relationship between the critic’s learned value function and the cost function of the differentiable MPC, providing a deeper understanding of the interplay between the critic’s value and the MPC cost functions. Finally, we validate our method in the drone racing task in various tracks, in both simulation and the real world. Our method achieves the same superhuman performance as state-of-the-art model-free RL, showcasing speeds of up to 21 m/s. We show that the proposed architecture can achieve real-time control performance, learn complex behaviors via trial and error, and retain the predictive properties of the MPC to better handle out of distribution behavior.
Supplementary Material
Video of the experiments: https://youtu.be/3epWs63bOGs
@@ I Introduction
Figure 1: Top: A block diagram of an actor-critic reinforcement learning architecture with a Multilayer Percepton (MLP) Bottom: A block diagram of the proposed approach. We combine the strength of actor-critic RL and the robustness of MPC by placing a differentiable MPC as the last module of the actor policy. At deployment time the commands for the environment are drawn from solving an MPC, which leverages the dynamics of the system and finds the optimal solution given the current state. We show that the proposed approach achieves better out-of-distribution behaviour and better robustness to changes in the dynamics. We also show that the predictions of the differentiable MPC can be used to improve the learning of the value function and the sample efficiency.
The animal brain’s exceptional ability to quickly learn and adjust to complex behaviors stands out as one of its most remarkable traits, which remains largely unattained by robotic systems. This has often been attributed to the brain’s ability to make both immediate and long-term predictions about the consequences of its actions, and plan accordingly [1, 2, 3]. In the field of robotics and control theory, model-based control has demonstrated a wide array of tasks with commendable reliability [4, 5]. In particular, Model Predictive Control (MPC) has achieved notable success across various domains such as the operation of industrial chemical plants [6], control of legged robots [7], and agile flight with drones [8, 9, 10, 11]. The effectiveness of MPC stems from its innate capability for online replanning. This enables it to make decisions that optimize a system’s future states over a specified short time horizon.
@@
Most recently, reinforcement-learning-based control has gained considerable traction, demonstrating exceptional performance in various domains, such as board games [17], atari games [18], complex video games [19, 20], and drone racing, both state-based [14, 21] and vision-based [22, 23]. In particular, the results in [14] demonstrate that model-free RL can achieve superior performance compared to optimal control techniques in drone racing. This success is largely due to RL’s ability to directly optimize a non-differentiable, nonlinear, sparse objective, eliminating the need for proxy objectives in the form of a predefined reference time trajectory or continuous 3D path. Additionally, RL learns a control policy through offline optimization, enabling efficient real-time control during deployment. Unlike trajectory optimization, it learns a feedback controller for real-time adaptability and directly optimizes task objectives, bypassing intermediate representations. Finally, RL can leverage domain randomization to cope with model uncertainty, allowing the discovery of more robust control responses. Unlike standard optimal control methods, RL optimizes policies directly through interactions with the environment. This allows RL to handle sparse reward functions (e.g., minimize number of crashes, minimize total energy consumption, minimize lap times, etc.). Additionally, RL can enhance robustness and prevent overfitting through domain randomization, which involves training the agent across a range of simulated environments with diverse settings. This flexibility and adaptability can lead to more scalable solutions, particularly in complex environments where MPC may struggle to provide optimal solutions [14].
However, RL architectures are not without their own set of challenges [15, 24]. Learning a model-free, end-to-end policy without explicitly leveraging prior knowledge in the training process, such as physics or dynamic models, results in the need to learn everything from data – often resulting in millions of interactions within a simulator. While the end-to-end paradigm is attractive, it often lacks in terms of generalizability and robustness to out-of-distribution scenarios. This has resulted in hesitancy in applying end-to-end learned architectures to safety-critical applications and has fostered the development of approaches that advocate for the introduction of safety in learned pipelines [24, 25, 26].
This article is an extension of a previous work [27]. In [27], we introduced a new architecture called Actor-Critic Model Predictive Control to bridge the gap between Reinforcement Learning and Model Predictive Control. This architecture equips the agent with a differentiable MPC [28], placed as the last module of the actor network, as shown in Fig. 1, that provides the system with online replanning capabilities and allows the policy to predict and optimize the short-term consequences of its actions.
@@
In [27], we demonstrated that the AC-MPC architecture exhibits more stable behavior in out-of-distribution scenarios, highlighting its robustness and generalizability compared to its neural-network-only counterpart. Lastly, in [27], we showcased the applicability of the proposed approach in real-world scenarios using an agile quadrotor platform, demonstrating the feasibility and effectiveness of the AC-MPC architecture in dynamic and challenging environments.
@@
  • We propose an enhancement to the AC-MPC algorithm by incorporating predictions into the critic’s training process. This improvement leverages the prior knowledge of the system through the MPC predictions and incorporates them in the learning process the value function, resulting in improved sample efficiency during training.
  • We conduct an empirical analysis that reveals a relationship between the Hessian of the learned value function and the learned cost functions of the differentiable MPC. This finding provides a deeper understanding of the interplay between the RL value and the MPC cost functions within the AC-MPC framework, offering valuable insights for future research.
  • We show that after optimizing hyperparameters, specifically exploration, AC-MPC is able to leverage the prior knowledge included in the dynamics of the system and achieve better training performance than AC-MLP (conventional PPO).
  • We show that our approach is more robust to dynamic parameter changes in comparison with a baseline, keeping high success rates for variations in mass, in XY inertia, and body rate limits. This improved robustness is attributed to the differentiable MPC’s access to the dynamics model.
  • We demonstrate through both simulation and real-world experiments that the proposed AC-MPC framework achieves performance on par with state-of-the-art model-free reinforcement learning approaches in the drone racing task, attaining speeds of up to 21 m/s on a quadrotor platform.
Overall, the AC-MPC architecture not only advances the integration of RL and MPC but also offers practical improvements in out of distribution behaviour, training efficiency, and real-world applicability, contributing to the development of more robust and efficient control systems.
@@ II Related Work
Another promising model-based direction for addressing complex control challenges is sampling-based MPC algorithms [55, 56, 57, 58], which are designed to handle intricate, non-differentiable cost functions and general nonlinear dynamics. These algorithms integrate system dynamics – either known or learned – into the Model Predictive Path Integral (MPPI) control framework, optimizing control in real-time. A key feature of sampling-based MPC is the ability to generate a large number of control samples on-the-fly, often leveraging parallel computation via Graphics Processing Units (GPUs). However, implementing sampling-based MPC on embedded systems poses significant challenges, as it tends to be both computationally demanding and memory-intensive. Reinforcement Learning (RL) in robotics. RL has risen as an attractive alternative to conventional controller design, achieving impressive performance that goes beyond model-based control in a variety of domains [17, 59, 19, 60, 61, 62]. RL optimizes a controller using sampled data and can handle nonlinear, nonconvex, non-differentiable and even sparse objectives, manifesting great flexibility in controller design. Compared to model-based optimal control, RL has a number of key advantages: most importantly, RL can optimize the performance objective of the task directly, removing the need for explicit intermediate representations, such as trajectories. Moreover, RL can solve complex tasks from raw, high-dimensional sensory input, without the need of a specific metric state [14]. However, RL is still far from reaching the level of robustness and generalisability of model-based control approaches when deployed in the real world, due to its brittleness when deployed in situations outside of the training distribution, and its lack of guarantees [24]. Model-free methods generally use black-box optimization methods and do not exploit the first order gradient through the dynamics, thus cannot leverage the full advantage of the prior knowledge.
@@ III Methodology
In this section we first present preliminaries about both Optimal Control and Reinforcement Learning, and then introduce our method.
@@ III-B Tracking Model Predictive Control
For tracking Model Predictive Control approaches [105, 106, 107], the objective $J(\bm{x})$ is to minimize a quadratic penalty on the error between the predicted states and inputs, and a given dynamically feasible reference $\bm{x}_{k,ref}$ and $\bm{u}_{k,ref}$. Consequently, the cost function $J(\bm{x})$ in problem (1) is substituted by:
@@
where $\Delta\bm{x}_{k}=\bm{x}_{k}-\bm{x}_{k,ref}$, $\Delta\bm{u}_{k}=\bm{u}_{k}-\bm{u}_{k,ref}$, and where $\bm{Q}\succeq 0$, $\bm{R}\succ 0$ and $\bm{P}\succeq 0$ are the state, input and final state weighting matrices. The norms of the form $\|\cdot\|_{A}^{2}$ represent the weighted Euclidean inner product $\|\bm{v}\|^{2}_{A}=\bm{v}^{T}A\bm{v}$.
This formulation relies on the fact that a feasible reference $\bm{x}_{k,ref}$, $\bm{u}_{k,ref}$ is accessible for the future time horizon $N$. Searching for these is often referred to as planning, which, depending on the application, can be both computationally intensive and complex, particularly in scenarios involving cluttered or dynamic environments [9, 108, 109]. In many cases, solving the planning problem becomes the bottleneck in real-time applications due to high computational requirements and the need for precise system models.
However, the tracking MPC formulation also relies on the assumption that a trajectory represents a well-posed, quadratic proxy objective for solving the end task. While this assumption holds for many practical use cases, it becomes limiting in situations where the system operates near its performance limits, such as aggressive maneuvers or high-speed tasks, where the trajectory may no longer serve as an effective proxy. In such cases, the tracking objective can result in suboptimal performance or even task failure [14].
@@ III-C Economic Model Predictive Control
On the other hand, economic MPC offers an alternative by directly optimizing the task objective. Unlike traditional tracking MPC, economic MPC allows for a more flexible design of the cost function, which can vary depending on the specific task [110, 111].
@@
While economic MPC offers advantages in terms of flexibility with respect to task design, it also presents several challenges that must be addressed for successful implementation. One of the primary caveats is that the stage cost $l(x_{k},u_{k})$ in most tasks tends to be sparse, non-differentiable and non-smooth, which complicates the formulation of Eq. (3). For example, in tasks with discontinuous behaviour – e.g., switching between different modes, dealing with contact forces, or dealing with agent crashes – it is generally difficult to find a cost function that is differentiable.
@@ III-D General Quadratic MPC formulation
In most MPC approaches there is a need of i) an explicit manual selection of a differentiable cost function that properly encodes the end task, and ii) hyperparameter tuning. As mentioned in section III-B, in the case of a standard tracking MPC this encoding is done through planning, by finding a dynamically feasible reference trajectory that translates the task into suitable cost function coefficients for every time step. For economic MPC (section III-C), recent research has been going in the direction of using data to approximate the cost function [112], or making it computationally tractable by transforming it to a quadratic program using the Hessian [113, 114].
However, these approaches present two main drawbacks: i) hand-crafting a dense, differentiable cost function can be difficult for a general task, and ii) even if this cost function is found, extra effort needs to be spent in fine tuning the hyperparameters for real-world deployment. More generally, optimization-based architectures such as MPC need to run in real-time when a deployment in the real world is desired. As a result, the underlying optimization problem is often simplified by approximating the original nonlinear formulation and converting it into a Quadratic Program (QP) suitable for the real-time iteration (RTI) [115, 116, 117] scheme to ensure computational tractability. A general quadratic cost function can be written as
@@
In this paper, we propose to directly search for the matrix coefficients of Eq. (4) – $Q_{k}$ and $p_{k}$ – using reinforcement learning and differentiable MPC [28]. This way we are able to encode the task as a general, non-differentiable reward function while at the same time maintaining computational tractability.
@@ III-E Actor-Critic Reinforcement Learning
To solve such problems, a common approach in Reinforcement Learning is to optimize directly the policy $\pi_{\theta}$ by using the policy gradient. The policy $\pi_{\theta}$ is parameterized by $\theta$, typically weights of a neural network. The goal is to adjust these parameters to maximize the expected return, defined in the infinite-horizon case as:
@@ III-F Actor-Critic Model Predictive Control
This paper proposes an Actor-Critic MPC controller architecture where the MPC is differentiable [28] and the cost function is learned end-to-end using RL. The MPC block is introduced as the differentiable module of the actor in an actor-critic pipeline which uses the proximal policy optimization (PPO) algorithm [119]. Our architecture is shown in Fig. 1. In contrast to previous work [120, 64] where the MPC is taken as a black-box controller and the gradient is sampled, in our case the gradient of the cost function with respect to the solution is analytically computed and propagated using a differentiable MPC [28]. Therefore, for every backward and forward pass of the actor network, we need to solve an optimization problem. Instead of resorting to task specific engineering of the cost function, we propose a neural cost map where the $Q_{k}$ and $p_{k}$ terms are the output of the neural network layers preceding the differentiable MPC block. This allows to encode the end task directly as a reward function, which is then trainable end-to-end using the PPO training scheme. The main benefit of this approach with respect to training a standard Multi Layer Perceptron (MLP) end-to-end is that the final module of the actor is a model-based MPC controller,
@@
and therefore it retains its generalizability and robustness properties. The model-based controller placed as the last module ensures that the commands are always feasible for the dynamics at hand, and that they respect the system constraints. This differentiable MPC supports control input constraints but not state constraints, which is why we add $\bm{u}_{k}\in\mathcal{U}$ in (1). To allow for exploration, during training the control inputs are sampled from a Gaussian distribution where the mean is the output of the MPC block, and the variance is controlled by the PPO algorithm. However, during deployment the output from the MPC is used directly on the system without further sampling, retaining all properties of a model-based controller.

### v7

@@ Abstract
An open research question in robotics is how to combine the benefits of model-free reinforcement learning (RL)—known for its strong task performance and flexibility in optimizing general reward formulations—with the robustness and online replanning capabilities of model predictive control (MPC). This paper provides an answer by introducing a new framework called Actor-Critic Model Predictive Control. The key idea is to embed a differentiable MPC within an actor-critic RL framework. This integration allows for short-term predictive optimization of control actions through MPC, while leveraging RL for end-to-end learning and exploration over longer horizons. Through various ablation studies, we expose the benefits of the proposed approach: it achieves better out-of-distribution behaviour, better robustness to changes in the dynamics and improved sample efficiency. Additionally, we conduct an empirical analysis that reveals a relationship between the critic’s learned value function and the cost function of the differentiable MPC, providing a deeper understanding of the interplay between the critic’s value and the MPC cost functions. Finally, we validate our method in a drone racing task on different tracks, in both simulation and the real world. Our method achieves the same superhuman performance as state-of-the-art model-free RL, showcasing speeds of up to 21 m/s. We show that the proposed architecture can achieve real-time control performance, learn complex behaviors via trial and error, and retain the predictive properties of the MPC to better handle out-of-distribution behavior.
@@ I Introduction
Figure 1: Top: A block diagram of an actor-critic reinforcement learning architecture with a Multilayer Perceptron (MLP) Bottom: A block diagram of the proposed approach. We combine the strength of actor-critic RL and the robustness of MPC by placing a differentiable MPC as the last module of the actor policy. At deployment time, the commands for the environment are drawn from solving an MPC, which leverages the system’s dynamics and finds the optimal solution given the current state. We show that the proposed approach achieves better out-of-distribution behavior and better robustness to changes in the dynamics. We also show that the predictions of the differentiable MPC can be used to improve the learning of the value function and the sample efficiency.
The animal brain’s exceptional ability to quickly learn and adjust to complex behaviors is one of its most remarkable traits, which remains unattained by robotic systems. This has often been attributed to the brain’s ability to make both immediate and long-term predictions about the consequences of its actions and plan accordingly [1, 2, 3]. In robotics and control theory, model-based control has demonstrated a wide array of tasks with commendable reliability [4, 5]. In particular, Model Predictive Control (MPC) has achieved notable success across various domains such as the operation of industrial chemical plants [6], control of legged robots [7], and agile flight with drones [8, 9, 10, 11]. The effectiveness of MPC stems from its innate capability for online replanning. This enables it to make decisions that optimize a system’s future states over a specified short time horizon.
@@
Most recently, reinforcement-learning-based control has gained considerable traction, demonstrating human-champion level or even superhuman performance in various domains, such as board games [17], atari games [18], complex video games [19, 20], and, in the physical world, in drone racing, both state-based [14] and vision-based [21]. In particular, the results in [14] demonstrate that model-free RL can achieve superior performance compared to optimal control techniques when pushing the limits in autonomous racing. This success is mainly due to the ability of RL to directly optimize a non-differentiable objective, eliminating the need for proxy objectives in the form of a predefined reference time trajectory or continuous 3D path. More generally, RL offers a flexible approach to control by directly optimizing a feedback policy through interactions with the environment, eliminating the need for intermediate representations. Unlike trajectory optimization or standard optimal control methods, RL can handle sparse reward signals (such as minimizing crashes, energy consumption, or lap times) and can improve robustness by leveraging domain randomization across diverse simulated conditions. This combination of offline learning, real-time adaptability, and robustness against uncertainty often makes RL more scalable and effective in complex scenarios where optimal control algorithms may struggle to yield real-time optimal solutions [14].
However, RL architectures are not without their own set of challenges [15, 22]. Learning a model-free, end-to-end policy without explicitly leveraging prior knowledge in the training process, such as physics or dynamic models, results in the need to learn everything from data, often resulting in millions of interactions within a simulator. While the end-to-end paradigm is attractive, it often lacks generalizability and robustness to out-of-distribution scenarios. This has resulted in hesitancy to apply end-to-end learned architectures to safety-critical applications and has fostered the development of approaches that advocate for the introduction of safety in learned pipelines [22, 23, 24].
This article is an extension of [25], which introduced an architecture called Actor-Critic Model Predictive Control to bridge the gap between Reinforcement Learning and Model Predictive Control. This architecture equips the agent with a differentiable MPC [26], placed after the last layer of the actor network, as shown in Fig. 1. This differentiable MPC module provides the system with online replanning capabilities and allows the policy to predict and optimize the short-term consequences of its actions.
@@
[25] demonstrated that the AC-MPC architecture exhibits more stable behavior in out-of-distribution scenarios, highlighting its robustness and generalizability compared to its neural-network-only counterpart. Lastly, [25] showcased the applicability of the proposed approach in real-world scenarios using an agile quadrotor platform, demonstrating the feasibility and effectiveness of the AC-MPC architecture in dynamic and challenging environments.
@@
  • It improves the AC-MPC algorithm by incorporating predictions into the critic’s training process. This improvement leverages the prior knowledge of the system through the MPC predictions and incorporates them in the learning process of the value function, resulting in improved sample efficiency during training.
  • It conducts an empirical analysis that reveals a relationship between the Hessian of the learned value function and the learned cost functions of the differentiable MPC. This finding provides a deeper understanding of the interplay between the RL value and the MPC cost functions within the AC-MPC framework, offering valuable insights for future research.
  • It shows that after optimizing hyperparameters, specifically exploration, AC-MPC can leverage the prior knowledge included in the dynamics of the system and achieve better training performance than AC-MLP.
  • It shows that AC-MPC is more robust to dynamic parameter changes than the AC-MLP baseline, keeping high success rates for variations in mass, in XY inertia, and body rate limits. This improved robustness is attributed to the differentiable MPC’s access to the dynamics model.
  • It demonstrates in both simulation and real-world experiments that the proposed AC-MPC achieves superhuman performance on par with state-of-the-art model-free reinforcement learning in a drone racing task, attaining speeds of up to 21 m/s on a quadrotor platform.
Overall, the AC-MPC architecture not only advances the integration of RL and MPC but also offers practical improvements in out-of-distribution behavior, training efficiency, and real-world applicability, contributing to the development of more robust and efficient control systems.
@@ II Related Work
Another promising model-based direction for addressing complex control challenges is sampling-based MPC algorithms [53, 54, 55, 56], which are designed to handle intricate, non-differentiable cost functions and general nonlinear dynamics. These algorithms integrate system dynamics – either known or learned – into the Model Predictive Path Integral (MPPI) control framework, optimizing control in real-time. A key feature of sampling-based MPC is the ability to generate a large number of control samples on-the-fly, often leveraging parallel computation via Graphics Processing Units (GPUs). However, implementing sampling-based MPC on embedded systems poses significant challenges, as it tends to be both computationally demanding and memory-intensive.
Reinforcement Learning (RL) in robotics. RL has risen as an attractive alternative to conventional controller design, achieving impressive performance that goes beyond model-based control in a variety of domains [17, 57, 19, 58, 59, 60]. RL optimizes a controller using sampled data and can handle nonlinear, nonconvex, non-differentiable and even sparse objectives, manifesting great flexibility in controller design. Compared to model-based optimal control, RL has a number of key advantages: most importantly, RL can optimize the performance objective of the task directly, removing the need for explicit intermediate representations, such as trajectories. Moreover, RL can solve complex tasks from raw, high-dimensional sensory input, without the need of a specific metric state [14]. However, RL is still far from reaching the level of robustness and generalisability of model-based control approaches when deployed in the real world, due to its brittleness when deployed in situations outside of the training distribution, and its lack of guarantees [22]. Model-free methods generally use black-box optimization methods and do not exploit the first-order gradient through the dynamics, thus cannot leverage the full advantage of the prior knowledge.
@@ III Methodology
In this section, we first present preliminaries about both Optimal Control and Reinforcement Learning, and then introduce our method.
@@ III-B Tracking Model Predictive Control
One of the primary uses of MPC is to control a dynamical system’s behavior such that it closely follows a pre-computed reference trajectory. This is known as trajectory tracking, and the corresponding MPC formulation is referred to as tracking MPC. In tracking MPC approaches [103, 104, 105], the objective $J(\bm{x})$ is to minimize a quadratic penalty on the error between the predicted states and inputs, and a given dynamically feasible reference $\bm{x}_{k,ref}$ and $\bm{u}_{k,ref}$. Consequently, the cost function $J(\bm{x})$ in problem (1) is substituted by:
@@
where $\Delta\bm{x}_{k}=\bm{x}_{k}-\bm{x}_{k,ref}$, $\Delta\bm{u}_{k}=\bm{u}_{k}-\bm{u}_{k,ref}$, and where $\bm{Q}\succeq 0$, $\bm{R}\succ 0$ and $\bm{P}\succeq 0$ are the state, input, and final state weighting matrices. The norms of the form $\|\cdot\|_{A}^{2}$ represent the weighted Euclidean inner product $\|\bm{v}\|^{2}_{A}=\bm{v}^{T}A\bm{v}$.
This formulation relies on the fact that a feasible reference $\bm{x}_{k,ref}$, $\bm{u}_{k,ref}$ is accessible for $N$ time steps in the horizon. Searching for this reference trajectory is often referred to as planning, which, depending on the application, can be both computationally intensive and complex, particularly in scenarios involving complex dynamics or cluttered environments [9, 106, 107]. In many cases, solving the planning problem becomes the bottleneck in real-time applications due to high computational requirements and the need for precise system models.
The tracking MPC formulation also relies on the assumption that a trajectory represents a well-posed, quadratic proxy objective for solving the end task. While this assumption holds for many practical use cases, it becomes limiting in situations where the system operates near its performance limits, such as aggressive maneuvers or high-speed tasks, where the trajectory may no longer serve as an effective proxy. In such cases, the tracking objective can result in suboptimal performance or even task failure [14].
@@ III-C Economic Model Predictive Control
To overcome these limitations, Economic MPC directly incorporates the task objective into the optimization problem, eliminating the need for a pre-defined reference trajectory. This fundamental difference allows for a more flexible cost function design, enabling the controller to be tailored to the specific economic objectives of the task [108, 109]:
@@
While economic MPC offers advantages in terms of flexibility with respect to task design, it also presents several challenges that must be addressed for successful implementation. One of the primary caveats is that the stage cost $l(x_{k},u_{k})$ in most tasks tends to be sparse and non-differentiable, which complicates the formulation of Eq. (3). For example, in tasks with discontinuous behavior (e.g., switching between different modes, dealing with contact forces, or dealing with agent crashes), it is generally difficult to find a cost function that is differentiable.
@@ III-D General Quadratic MPC formulation
In most MPC approaches, there is a need of i) an explicit manual selection of a differentiable cost function that properly encodes the end task, and ii) hyperparameter tuning. As mentioned in section III-B, in the case of a standard tracking MPC this encoding is done through planning, by finding a dynamically feasible reference trajectory that translates the task into suitable cost function coefficients for every time step. For economic MPC (section III-C), recent research has been going in the direction of using data to approximate the cost function [110], or making it computationally tractable by transforming it to a quadratic program using the Hessian [111, 112].
However, these approaches present two main drawbacks: i) hand-crafting a dense, differentiable cost function can be difficult for a general task, and ii) even if this cost function is found, extra effort needs to be spent in fine tuning the hyperparameters for real-world deployment. More generally, optimization-based architectures such as MPC need to run in real-time when deployment in the real world is desired. As a result, the underlying optimization problem is often simplified by approximating the original nonlinear formulation and converting it into a Quadratic Program (QP) suitable for the real-time iteration (RTI) [113, 114, 115] scheme to ensure computational tractability. A general quadratic cost function can be written as
@@
In this paper, we propose to directly search for the matrix coefficients of Eq. (4) (i.e., $Q_{k}$ and $p_{k}$) using reinforcement learning and differentiable MPC [26]. This way, we can encode the task as a general, non-differentiable reward function while at the same time maintaining computational tractability.
@@ III-E Actor-Critic Reinforcement Learning
To solve such problems, a common approach in Reinforcement Learning is to optimize directly the policy $\pi_{\theta}$ by using the policy gradient. The policy $\pi_{\theta}$ is parameterized by $\theta$, typically the weights of a neural network. The goal is to adjust these parameters to maximize the expected return, defined in the infinite-horizon case as:
@@ III-F Actor-Critic Model Predictive Control
This paper proposes an Actor-Critic MPC controller architecture where the MPC is differentiable [26] and the cost function is learned end-to-end using RL. The MPC block is introduced as the differentiable module of the actor in an actor-critic pipeline that uses the proximal policy optimization (PPO) algorithm [117]. Our architecture is shown in Fig. 1. In contrast to previous work [118, 62], where the MPC is taken as a black-box controller and the gradient is sampled, in our case the gradient of the cost function with respect to the solution is analytically computed and propagated using a differentiable MPC [26]. Therefore, for every backward and forward pass of the actor network, we need to solve an optimization problem. Instead of resorting to task-specific engineering of the cost function, we propose a neural cost map where the $Q_{k}$ and $p_{k}$ terms are the output of the neural network layers preceding the differentiable MPC block. This allows us to encode the end task directly as a reward function, which is then trainable end-to-end using the PPO training scheme. The main benefit of this approach with respect to training a standard Multi-Layer Perceptron (MLP) end-to-end is that the final module of the actor is a model-based MPC controller,
@@
and therefore it retains its generalizability and robustness properties. The model-based controller placed after the last layer of the actor network ensures that the commands are always feasible for the dynamics at hand, and that they respect the system constraints. This differentiable MPC supports control input constraints but not state constraints, which is why we add $\bm{u}_{k}\in\mathcal{U}$ in (1). To allow for exploration, during training the control inputs are sampled from a Gaussian distribution where the mean is the output of the MPC block, and the variance is controlled by the PPO algorithm. However, during deployment the output from the MPC is used directly on the system without further sampling, retaining all properties of a model-based controller.